OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of -1/5

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] Slope=m x1,y1

pooja195 (pooja195):

Can you sub in the info?

OpenStudy (jackellyn):

y=mx+b is the slope intercept form m is slope and b is the y intercept. y=-1/5x+b Use the point (-3,7) to find the y intercept, b. Or use the formula @pooja195 said.